OpenStudy (anonymous):

Vectors Projectile Motion: A tiger leaps horizontally from a 7.0m high rock with a speed of 3.0m/s.How far from the base of the rock will she land?

OpenStudy (anonymous):

I know I have to find the final velocity of the tiger at the rock, I just want to make sure I've set up everything correctly. Since the tiger leaps horizontally, gravity is negative right?

OpenStudy (ash2326):

Let's find the time it takes for the tiger to fall under gravity from height 7 meters, its initial vertical velocity is zero \[S=ut+\frac{1}{2} at^2\] u=initial velocity=0 S=7 meters a=9.8 m/s^2 \[7=\frac{1}{2}\times 9.8 \times t^2\] \[\frac{14}{9.8}=t^2\] we get \[t=1.195 seconds\] In this time it'll cover horizontal distance say D \[D=3 m/s \times 1.195 s=3.585 m\]

OpenStudy (anonymous):

no.because gravity will be negative if the tiger would leap upward direction

OpenStudy (ash2326):

It depends which direction we assume negative, if I assume downward direction as positive then gravity is positive, motion in downward direction is positive. If we assume upward direction as positive then gravity will be negative. Here I've assumed downward direction as positive

OpenStudy (ash2326):

@JerJason Did you understand?

OpenStudy (anonymous):

Yes, I understood how you went about getting the answers, I was just confused about how to tell whether gravity was negative or position but you both explained it for me. Thanks.

OpenStudy (ash2326):

Glad to help:D